Solve: log(base 2) x - 2log(base2) (x-1) = 1
Answer by lwsshak3(11628)

Solve: log(base 2) x - 2log(base2) (x-1) = 1
log2(x)-2log2(x-1)=1
log2[x/(x-1)^2]=1
convert to exponential form
2^1=[x/(x-1)^2=2
2(x-1)^2]=x
2(x^2-2x+1)=x
2x^2-4x+2=x
2x^2-5x+2=0
(2x-1)(x-2)=0
x=1/2 (reject,(x-1)>0)
or
x=2